6) A piece of metal at 299oC is immersed in 361 ml of water at 23.2oC. The water temperature rises to 63.3oC. SH for the metal is 0.361 J/g-K

a. What is the mass of the piece of steel?

b. How precise is this method for measuring mass?

c. Explain b
Answer #1

a. Heat released by metal = heat absorbed by water

Mass of steel * Specific heat of steel * Temperature change = mass of water * Specific heat of water * Temperature change

Mass of water = Volume * Density = 361 * 1 = 361 g

m * 0.361 * (299 - 63.3) = 361 * 4.184 * (63.3 - 23.2)

Solving, m = 711.83 g

b. This is not a very accurate measure of mass

c. The reason lies in the fact that we cannot fully guarantee that all of the released heat goes into the water meaning the measured mass is actually different from the actual mass.